OpenStudy (anonymous):

Name 3 ordered pairs which belong to the relation 2x+3y=6 ~~How do you do this question?

OpenStudy (mathmate):

Set a value for y and solve for x. That's one ordered pair. Repeat for other values of y to get two other pairs. Example: y=0, x=6/2=3, so (3,0) is an ordered pair belonging to the relation.

OpenStudy (anonymous):

so you just pick random numbers for x and y and solve them?

OpenStudy (mathmate):

Yes, since the domain of the relations is any real number.

OpenStudy (anonymous):

if you pick 0 for x would it be 2(0)+3y=6 and then pick another number for y?

OpenStudy (mathmate):

Yes, you can pick a number for x as well.

jimthompson5910 (jim_thompson5910):

if you pick 0 for x, you plug it in to get 2(0)+3y=6 0+3y = 6 3y = 6 Now solve for y to get 3y = 6 3y/3 = 6/3 y = 2 So when x = 0, y = 2 So we have one ordered pair (0, 2)

OpenStudy (mathmate):

If you pick x=0, then you solve for y=6/2=2,so (0,2) is another pair.